OpenStudy (anonymous):

A concrete cube of side 0.60 m and uniform density 2.0 × 103 kg m–3 is lifted 5.0 m vertically by a crane. What is the change in potential energy of the cube?

OpenStudy (mos1635):

formula for potential energy??

OpenStudy (anonymous):

i just need an answer if however i require the formula

OpenStudy (mos1635):

????

OpenStudy (anonymous):

these are the answers: A 2.2 kJ B 21 kJ C 59 kJ D 450 kJ

OpenStudy (anonymous):

its B but i wanna know how to get the answer

OpenStudy (mos1635):

what is the formula for potential energy??

OpenStudy (anonymous):

i know that its equal to mgh m for mass, g equals to 9.81 and h for height but i still can solve it; density is related as: mass over volume and well im not sure how to solve it

OpenStudy (mos1635):



OpenStudy (anonymous):

0.6 X3 as the sides of the cube?

OpenStudy (mos1635):

0.6^3=0.6*0.6*0.6=0.216

OpenStudy (anonymous):

ok and next?

OpenStudy (mos1635):

find mass

OpenStudy (anonymous):

we have the density so i guess we need the volume and we find the mass

OpenStudy (mos1635):

\[d=\frac{ m }{ V }\]

OpenStudy (mos1635):

but you have allready find volume

OpenStudy (anonymous):

oh yes its 0.216

OpenStudy (mos1635):

\[V _{cube}=side ^{3}=a ^{3}=0.6^{3}=0.216\]

OpenStudy (anonymous):

ok thanks just found it thanks for your help :)